# A Test of Substellar Evolutionary Models with High-Precision Ages from Asteroseismology and Gyrochronology for the Benchmark System HR 7672AB

### v1

@@ 4 Testing Brown Dwarf Cooling Models
References — An25: 2025ApJS..280...61A; Balm23: 2023ApJ...956...99B; Balm24: 2024AJ....167...64B; Balm25: 2025AJ....169...30B; Bard21: 2021ApJ...922L..43B; Bard25: 2025ApJ...988L..18B; Bell15: 2015MNRAS.454..593B; Belo18: 2018MNRAS.478..611B; Bill10: 2010ApJ...720L..82B; Bona22: 2022MNRAS.513.5588B; Bouc16: 2016A&A...585A..46B; Bowl18: 2018AJ....155..159B; Bran15: 2015ApJ...807...58B; Bran19: 2019AJ....158..140B; Bran20: 2020AJ....160..196B; Bran21a: 2021ApJ...915L..16B; Bran21b: 2021AJ....162..301B; Bran21c: 2021AJ....161..179B; Ceva25: 2025A&A...701A..78C; Chee18b: 2018A&A...614A..16C; Chen22: 2022AJ....163..288C; Chil17: 2017AJ....153..182C; Crep15: 2015ApJ...798L..43C; Crep16: 2016ApJ...831..136C; Curr20: 2020ApJ...904L..25C; Curr23: 2023Sci...380..198C; DeRo23: 2023A&A...672A..94D; Delo17b: 2017A&A...608A..79D; Dupu09a: 2009ApJ...692..729D; Dupu12: 2012ApJS..201...19D; Dupu14: 2014ApJ...790..133D; Dupu17: 2017ApJS..231...15D; ElMo25: 2025ApJ...981...20E; Fran22: 2022AJ....163...50F; Fran23: 2023ApJ...950L..19F; Fran23b: 2023AJ....165...39F; Fran23c: 2023AJ....165..246F; Goto02: 2002ApJ...567L..59G; Helm18: 2018Natur.563...85H; Henr90: 1990ApJ...354L..29H; Hink23: 2023A&A...671L...5H; Irel08: 2008ApJ...678..463I; Kamm24: 2024AJ....168...51K; Kamm25: 2025arXiv251008691K; King10: 2010A&A...510A..99K; Kirk00: 2000AJ....120..447K; Kono10: 2010ApJ...711.1087K; Kuzu22: 2022ApJ...934L..18K; Lagr10: 2010Sci...329...57L; Lagr19: 2019NatAs...3.1135L; Lane01: 2001ApJ...560..390L; Li23: 2023MNRAS.522.5622L; Li24: 2024MNRAS.533.3501L; Li26: this work; Liu02b: 2002ApJ...571..519L; Mair20a: 2020A&A...633L...2M; Mair20b: 2020A&A...639A..47M; Mair24: 2024A&A...691A.263M; Maro10b: 2010Natur.468.1080M; Matt24: 2024Natur.633..789M; Mesa20: 2020MNRAS.495.4279M; Mesa23: 2023A&A...672A..93M; Mesh15b: 2015MNRAS.453.2378M; Mill17: 2017A&A...597L...2M; Naka95b: 1995Natur.378..463N; Nase24: 2024A&A...687A.298N; Nowa20: 2020A&A...642L...2N; Peca16: 2016MNRAS.461..794P; Pere19: 2019A&A...631A.107P; Pott02: 2002ApJ...567L.133P; Prav05: 2005ApJ...630..528P; Rick19: 2019A&A...625A..71R; Rick20: 2020A&A...635A.203R; Rick24: 2024A&A...684A..88R; Scho03: 2003A&A...398L..29S; Skru87b: 1987BAAS...19.1128S; Thal09: 2009ApJ...707L.123T; Tobi24: 2024AJ....167..205T; Viga16: 2016A&A...587A..55V; Wagn20: 2020ApJ...902L...6W; Ward21: 2021AJ....161....5W; Wint24: 2024A&A...688A..44W; Wint25: 2025A&A...700A...4W; Xuan24: 2024Natur.634.1070X; Zurl16: 2016A&A...587A..57Z
@@
4.1 Consistency tests among mass, age, and luminosity
Independent measurements of mass, age, and luminosity allow any two quantities to predict the third using an evolutionary model, enabling a direct comparison between the model-predicted and observed values for the third quantity. We performed this test using the rejection-sampling procedure described in Dupuy2017 and Dupuy2023. To obtain the model-predicted distribution of $\log L$ based on mass and age, we drew normally distributed random values for mass and age according to their observational uncertainties and calculated the corresponding luminosity for each mass-age pair through bicubic interpolation on the model grids.
@@ 4.1 Consistency tests among mass, age, and luminosity
From these comparisons, we found that the Burrows1997 models predict the lowest luminosities and a clearly substellar nature for HR 7672B. The ATMO2020 model lies near the transition between the bottom of the main sequence and the brown dwarf regime, while its successor CBPD2023 shows the best agreement with our measurements and remains below the threshold for sustained hydrogen fusion. The three models that include cloud treatment through the L/T transition (BHAC2015, SM2008, and Diamondback) predict slightly higher luminosities than observed, indicating a stellar outcome for HR 7672B at the low-mass end of the main sequence given its measured dynamical mass. These models suggest that the object would reach the zero-age main sequence in approximately 1–2 Gyr, producing the long tail in the age distribution in Figure 9 and the flattening in the cooling tracks in Figure 7. The implication for these three models is that their predicted stellar/substellar boundary is too low in mass.
@@
Finally, we consider which observed quantities could most improve model discrimination (Figure 9). The age uncertainty (18%) is relatively large, and further improvements in the host star age will be challenging, given the substantial observing time needed to measure the current asteroseismic age. Moreover, an improved age is unlikely to enhance model separation, as the system lies in a region where models predict similar luminosities. The narrow substellar mass range already provides some discriminatory power, and upcoming data from Gaia DR4 can further refine the mass and improve model differentiation. A reduced luminosity uncertainty (currently 10%) would likely provide the strongest improvement, as it could distinguish between the higher-mass substellar boundary predicted by ATMO2020 and the lower-mass boundary suggested by models such as SM2008, which perform well across the L/T transition (2015ApJ...805...56D; Dupuy2017; Chenmh2022).
4.2 Goodness-of-fit test for mass, age, and luminosity
@@
We determined the distance $D^{2}$ between the observed point and the nearest point on each model manifold. The distance $D^{2}$ follows a chi-squared distribution with 1 degree of freedom, corresponding to the single direction normal to the two-dimensional manifold in three-dimensional parameter space (3–2). We calculated the one-sided tail probability $p$, defined as the probability that a random draw from this distribution exceeds the observed $D^{2}$. The probability $p$ is then converted to an equivalent $\sigma$ value using the inverse survival function of a normal distribution. The results are reported in Table 9.
@@
5 Near-Simultaneous Asteroseismology with TESS
Because oscillation frequencies are detected in both the TESS and KPF datasets, HR 7672 offers an opportunity to constrain photometry-to-RV amplitude ratios. These ratios have previously been measured in only a few stars with simultaneous data, including Procyon (Huber2011-procyon), HD 35833 (Gupta2022), and $\beta$ Aql (Kjeldsen2025). We follow the approach of Kjeldsen2008 to measure oscillation amplitudes. Each power density spectrum — defined as the power spectrum multiplied by the effective observing duration — was smoothed with a Gaussian kernel with a full width at half-maximum of $4{\Delta\nu}{}$ to produce a single power excess hump. The smoothed spectrum was then multiplied by ${\Delta\nu}{}/c$ to estimate the power per radial order, where $c$ accounts for the effective number of modes per order (Ballot2011). Given the star’s similarity to the Sun, we adopt solar values calculated by Kjeldsen2008: $c=4.09$ for velocity data and $c=2.91$ for TESS photometry. Figure 13 shows the resulting spectra. The square root of the maximum near $\nu_{\text{max}}$ yields the oscillation amplitude. Uncertainties in the oscillation amplitudes were estimated via bootstrapping: in each realization, the original spectrum was multiplied by a noise spectrum, and the oscillation amplitude was remeasured. The spread in amplitude estimates across all realizations was adopted as the uncertainty.
For the KPF data, we measured a RV amplitude of $16.9\pm 0.6$ cm s-1. In the TESS data, oscillations are not clearly visible as a distinct power excess due to dominant contributions from granulation and white noise. We therefore derive an upper limit of $2.8\pm 0.7$ ppm for the photometric amplitude in the TESS bandpass. Applying a bolometric correction factor of 1.2 (Lund2019), the corresponding bolometric amplitude is $3.4\pm 0.9$ ppm. These values yield an upper limit on the bolometric photometry-to-RV amplitude ratio of $20.1\pm 7.6$ ppm (m s-1)-1. Both the amplitudes and their ratio are similar to the solar values. For comparison, the solar oscillation amplitude is approximately 18.7 cm s-1 in RV (Kjeldsen2008) and 3.6 ppm in photometry (Michel2009), corresponding to a ratio of about 19.2 ppm (m s-1)-1.
Further, both the amplitudes and their ratio are broadly in line with predictions from 3D hydrodynamic simulations (zhouyx2021). Repeating the methods of zhouyx2021, we obtained a simulated solar amplitude ratio between bolometric intensity and fluid velocity of 20.12 ppm (m s-1)-1. For a hotter main-sequence model whose basic stellar parameters are close to those of HR 7672A (${T_{\text{eff}}}{}=6000$ K and ${\log g}{}=4.5$ dex), the amplitude ratio is slightly lower at 19.27 ppm (m s-1)-1. The trend predicted from simulations is also consistent with the empirical amplitude scaling relation of kb95. Assuming the conversion factor between fluid velocity and RV is invariant from the Sun to HR 7672A, which is reasonable given the similarity between the two stars, the expectation is that the actual photometry-to-RV amplitude ratio of the latter is slightly smaller than 19.2 ppm (m s-1)-1, in agreement with our measured upper limit.
Oscillation amplitudes are expected to be lower in stars with higher magnetic activity (Campante2014; Sayeed2025). Sayeed2025 quantified this dependency using the chromospheric activity levels derived from Ca II H&K line emission, using the logarithmic flux ratio between the emission flux and the bolometric flux, $\log R^{\prime}_{\rm HK}$. BoroSaikia2018 reported $\log R^{\prime}_{\rm HK}$ values between 4.68-4.68 and 4.74-4.74 for this star; for comparison, the solar value is around 4.94-4.94 (Egeland2017). According to Equation 10 of Sayeed2025, this corresponds to an amplitude ratio between 0.81 and 0.85 relative to the Sun. Although our measurement is broadly consistent with this prediction, it does not provide conclusive evidence for a reduced amplitude in this star due to the large uncertainties. The photometric-to-RV amplitude ratio, however, is likely less affected by magnetic activities, as it depends on the stellar structure near the photosphere to a first order (zhouyx2021). By comparing hydrodynamics and magnetohydrodynamics simulations of stellar atmospheres, 2022A&A...663A.166B show that changes in near-surface density, temperature, and pressure stratification caused by the small-scale dynamo are less than 1% for G-type stars.
6 Conclusions
@@ Acknowledgements
The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has benefited from The UltracoolSheet at http://bit.ly/UltracoolSheet, maintained by Will Best, Trent Dupuy, Michael Liu, Aniket Sanghi, Rob Siverd, and Zhoujian Zhang, and developed from compilations by 2012ApJS..201...19D; 2013Sci...341.1492D; 2014ApJ...792..119D; 2016ApJ...833...96L; 2018ApJS..234....1B; 2021AJ....161...42B; 2023ApJ...959...63S; 2023AJ....166..103S. We thank Will Balmer, Brendan Bowler, Kyle Franson, and Jerry Xuan for discussions about the census of substellar companions with dynamical masses. Y.L. acknowledges support from the National Aeronautics and Space Administration (80NSSC25K7904) and the Beatrice Watson Parrent Fellowship. M.L. acknowledges past support from the Parrent Fellowship 24 years ago that enabled the discovery of HR 7672B and current support from the Gordon and Betty Moore Foundation through grant GBMF8550. T.S.M. acknowledges support from NASA grant 80NSSC25K7898. T.R.B. acknowledges support from the Australian Research Council (FL220100117). T.L.C. is supported by Fundação para a Ciência e a Tecnologia (FCT) in the form of a work contract (2023.08117.CEECIND/CP2839/CT0004). J.R.L. and M.S.L. acknowledge support from VILLUM FONDEN (42101) and The Independent Research Fund Denmark’s Inge Lehmann program (1131-00014B).
@@ Appendix A Dynamical Orbit Fit
Figures 14 and 15 present the best-fitting dynamical orbit model to the astrometric and long-baseline radial velocity data. The corresponding parameter estimates are listed in Table 10.

### v2

@@ IV Testing Brown Dwarf Cooling Models
References — An25: Q. An et al. (2025b); Balm23: W. O. Balmer et al. (2023); Balm24: W. O. Balmer et al. (2024); Balm25: W. O. Balmer et al. (2025); Bard21: D. C. Bardalez Gagliuffi et al. (2021); Bard25: D. C. Bardalez Gagliuffi et al. (2025); Bell15: C. P. M. Bell et al. (2015); Belo18: V. Belokurov et al. (2018); Bill10: B. A. Biller et al. (2010); Bona22: M. Bonavita et al. (2022); Bouc16: F. Bouchy et al. (2016); Bowl18: B. P. Bowler et al. (2018); Bran15: T. D. Brandt & C. X. Huang (2015); Bran19: T. D. Brandt et al. (2019b); Bran20: T. D. Brandt et al. (2020); Bran21a: G. M. Brandt et al. (2021b); Bran21b: G. M. Brandt et al. (2021c); Bran21c: G. M. Brandt et al. (2021a); Ceva25: W. Ceva et al. (2025); Chee18b: A. Cheetham et al. (2018); Chen22: M. Chen et al. (2022a); Chil17: J. Chilcote et al. (2017); Crep15: J. R. Crepp et al. (2015); Crep16: J. R. Crepp et al. (2016); Curr20: T. Currie et al. (2020); Curr23: T. Currie et al. (2023); Curr25: T. Currie et al. (2025); DeRo23: R. J. De Rosa et al. (2023); Delo17b: P. Delorme et al. (2017); Dupu09a: T. J. Dupuy et al. (2009); Dupu12: T. J. Dupuy & M. C. Liu (2012); Dupu14: T. J. Dupuy et al. (2014); Dupu17: T. J. Dupuy & M. C. Liu (2017b); ElMo25: M. El Morsy et al. (2025); Fran22: K. Franson et al. (2022); Fran23: K. Franson et al. (2023a); Fran23b: K. Franson et al. (2023b); Fran23c: K. Franson & B. P. Bowler (2023); Goto02: M. Goto et al. (2002); Helm18: A. Helmi et al. (2018); Henr90: T. J. Henry & J. D. Kirkpatrick (1990); Hink23: S. Hinkley et al. (2023); Irel08: M. J. Ireland et al. (2008); Kamm24: J. Kammerer et al. (2024); Kamm25: J. Kammerer et al. (2025); King10: R. R. King et al. (2010); Kirk00: J. D. Kirkpatrick et al. (2000); Kono10: Q. M. Konopacky et al. (2010); Kuzu22: M. Kuzuhara et al. (2022); Lagr10: A. Lagrange et al. (2010); Lagr19: A.-M. Lagrange et al. (2019); Lane01: B. F. Lane et al. (2001); Li23: Y. Li et al. (2023b); Li24: Y. Li et al. (2024); Li26: this work; Liu02b: M. C. Liu et al. (2002b); Mair20a: A.-L. Maire et al. (2020a); Mair20b: A.-L. Maire et al. (2020b); Mair24: A.-L. Maire et al. (2024); Maro10b: C. Marois et al. (2010); Matt24: E. C. Matthews et al. (2024); Mesa20: D. Mesa et al. (2020); Mesa23: D. Mesa et al. (2023); Mesh15b: T. Meshkat et al. (2015); Mill17: J. Milli et al. (2017); Naka95b: T. Nakajima et al. (1995); Nase24: E. Nasedkin et al. (2024); Nowa20: M. Nowak et al. (2020); Peca16: M. J. Pecaut & E. E. Mamajek (2016); Pere19: S. Peretti et al. (2019); Pott02: D. Potter et al. (2002); Prav05: S. H. Pravdo et al. (2005); Rick19: E. L. Rickman et al. (2019); Rick20: E. L. Rickman et al. (2020); Rick24: E. L. Rickman et al. (2024); Scho03: R.-D. Scholz et al. (2003); Skru87b: M. F. Skrutskie et al. (1987); Thal09: C. Thalmann et al. (2009); Tobi24: T. L. Tobin et al. (2024); Viga16: A. Vigan et al. (2016); Wagn20: K. Wagner et al. (2020); Ward21: K. Ward-Duong et al. (2021); Wint24: T. O. Winterhalder et al. (2024); Wint25: T. O. Winterhalder et al. (2025); Xuan24: J. W. Xuan et al. (2024); Zurl16: A. Zurlo et al. (2016)
@@
IV.1 Consistency tests among mass, age, and luminosity
Independent measurements of mass, age, and luminosity allow any two quantities to predict the third using an evolutionary model, enabling a direct comparison between the model-predicted and observed values for the third quantity. We performed this test using the rejection-sampling procedure described in T. J. Dupuy & M. C. Liu (2017a) and T. J. Dupuy et al. (2023). To obtain the model-predicted distribution of $\log L$ based on mass and age, we drew normally distributed random values for mass and age according to their observational uncertainties and calculated the corresponding luminosity for each mass-age pair through bicubic interpolation on the model grids.
@@ IV.1 Consistency tests among mass, age, and luminosity
Figure 13: The luminosity differences between 8 and 10 Gyr, $\Delta\log(L/{\text{L}_{\odot}})$, shown as a function of mass. The measured stellar/substellar boundaries are shown in circles.
@@
Table 10: Predicted stellar/substellar boundary from each evolutionary model. For reference, HR 7672B’s mass is $75.39\pm 0.67$ $\text{M}_{\text{Jup}}$.
Model Boundary Mass ($\text{M}_{\text{Jup}}$)
Diamondback $69.2$
SM2008 $71.8$
BHAC2015 $72.3$
ATMO2020 $77.7$
CBPD2023 $78.6$
Burrows1997 $82.0$
From these comparisons, we found that the Burrows1997 models predict the lowest luminosities and a clearly substellar nature for HR 7672B. The ATMO2020 model lies near the transition between the bottom of the main sequence and the brown dwarf regime, while its successor CBPD2023 shows the best agreement with our measurements and remains below the threshold for sustained hydrogen fusion. The three models that include cloud treatment through the L/T transition (BHAC2015, SM2008, and Diamondback) predict slightly higher luminosities than observed, indicating a stellar outcome for HR 7672B at the low-mass end of the main sequence given its measured dynamical mass. These models suggest that the object would reach the zero-age main sequence in approximately 1–2 Gyr, producing the long tail in the age distribution in Figure 14 and the flattening in the cooling tracks in Figure 11. The implication for these three models is that their predicted stellar/substellar boundary is too low in mass. For reference, we list the stellar/substellar boundary in Table 10. To compute these values, we examined the brown dwarf cooling sequences and evaluated the change in $\log L$ between 8 and 10 Gyr. We denote the stellar/substellar boundary as the mass above which the $\log L$ change remains below 0.002 dex. Figure 13 illustrates this procedure.
@@
Finally, we consider which observed quantities could most improve model discrimination (Figure 14). The age uncertainty (18%) is relatively large, and further improvements in the host star age will be challenging, given the substantial observing time needed to measure the current asteroseismic age. Moreover, an improved age is unlikely to enhance model separation, as the system lies in a region where models predict similar luminosities. The narrow substellar mass range already provides some discriminatory power, and upcoming data from Gaia DR4 can further refine the mass and improve model differentiation. A reduced luminosity uncertainty (currently 10%) would likely provide the strongest improvement, as it could distinguish between the higher-mass substellar boundary predicted by ATMO2020 and the lower-mass boundary suggested by models such as SM2008, which perform well across the L/T transition (T. J. Dupuy et al., 2015; T. J. Dupuy & M. C. Liu, 2017a; M. Chen et al., 2022b).
IV.2 Goodness-of-fit test for mass, age, and luminosity
@@
We determined the distance $D^{2}$ between the observed point and the nearest point on each model manifold. The distance $D^{2}$ follows a chi-squared distribution with 1 degree of freedom, corresponding to the single direction normal to the two-dimensional manifold in three-dimensional parameter space (3–2). We calculated the two-sided tail probability $p$, defined as the probability that a random draw from this distribution exceeds the observed $D^{2}$. The probability $p$ is then converted to an equivalent $\sigma$ value using the inverse survival function of a normal distribution. The results are reported in Table 9.
@@
V Conclusions
@@ Acknowledgements
The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has benefited from The UltracoolSheet at http://bit.ly/UltracoolSheet, maintained by Will Best, Trent Dupuy, Michael Liu, Aniket Sanghi, Rob Siverd, and Zhoujian Zhang, and developed from compilations by T. J. Dupuy & M. C. Liu (2012); T. J. Dupuy & A. L. Kraus (2013); N. R. Deacon et al. (2014); M. C. Liu et al. (2016); W. M. J. Best et al. (2018, 2021); A. Sanghi et al. (2023); A. C. Schneider et al. (2023). We thank Will Balmer, Brendan Bowler, Kyle Franson, and Jerry Xuan for discussions about the census of substellar companions with dynamical masses, and Adam Burrows for comments on the manuscript. Y.L. acknowledges support from the National Aeronautics and Space Administration (80NSSC25K7904) and the Beatrice Watson Parrent Fellowship. M.L. acknowledges past support from the Parrent Fellowship 24 years ago that enabled the discovery of HR 7672B and current support from the Gordon and Betty Moore Foundation through grant GBMF8550. T.S.M. acknowledges support from NASA grant 80NSSC25K7898. T.R.B. acknowledges support from the Australian Research Council (FL220100117). T.L.C. is supported by Fundação para a Ciência e a Tecnologia (FCT) in the form of a work contract (2023.08117.CEECIND/CP2839/CT0004). J.R.L. and M.S.L. acknowledge support from VILLUM FONDEN (42101) and The Independent Research Fund Denmark’s Inge Lehmann program (1131-00014B).
@@ Appendix A Dynamical Orbit Fit
Figures 15 and 16 present the best-fitting dynamical orbit model to the astrometric and long-baseline radial velocity data. The corresponding parameter estimates are listed in Table 11.